A sewage lift station has tank (3000 L) that is stagnant for 8 h during low flows each night. The sewage contains 10 mg/L SO42--S. Sulfate (SO42-) is converted to sulfide (S2-), which is then converted to toxic H2S gas. The reaction rate of sulfate-sulfur to sulfide-sulfur can be described by a second order decay, k = 0.01 L/(mg∙h). How long (h) will it take for half of the sulfate-sulfur in the lift station to be converted to sulfide-sulfur?
Answer #1

for second order reaction, T= CAOXA/-rA

-rA= KCA2= KCAO2*(1-XA)2=

T= CAOXA/KCAO2*(1-XA)2=

K is the rate constant = 0.01 L/mg.h

XA=0.5 , CAO=10mg/L

T= (XA/(1-XA)2 * 1/KCAO = (0.5/0.25)*(1/0.01*10)= 2 hrs

the time taken = 2 hrs